OpenStudy (anonymous):

Make sure you read the directions for each part carefully. This system of equations will be used in all three parts of this question. 8x-2y=6 3x+y=4 Part 1: Solve the system by graphing. Draw the graphs on paper and indicate their intersection point. Type the ordered pair solution here as your answer to Part 1. Part 2: Solve the system using the substitution method. Show all work here and indicate the solution for the system as an ordered pair. Part 3: Solve the system using the addition method. Show all work here and indicate the solution for the system as an ordered pair.

OpenStudy (anonymous):

to solve by graphing first get it into y = mx + b format

OpenStudy (anonymous):

i dont understand this problem at all

OpenStudy (***[isuru]***):

hi... let's start this from the bottom.. do u know how to plot a linear function ? i mean a graph in y = mx + b format

OpenStudy (anonymous):

ya but i dont know the point of intersection

OpenStudy (***[isuru]***):

do u mean that u don't know the way u get an answer from graph ??

OpenStudy (***[isuru]***):

here's how... let say u have to solve a system of equation... like the one below using graphs y = ax + b 2y = -dx +3 then first u must draw them on a graph... 

OpenStudy (***[isuru]***):

then...